A mutual fund manager has a $20 million portfolio with a beta of 2.00. The risk-free rate is 6.75%, and the market risk premium is 5.0%. The manager expects to receive an additional $5 million, which she plans to invest in a number of stocks. After investing the additional funds, she wants the fund's required return to be 18%. What should be the average beta of the new stocks added to the portfolio? Do not round intermediate calculations. Round your answer to two decimal places. Enter a negative answer with a minus sign.
Answer #1

Expected Portfolio Return = 18%, Risk-Free Rate = Rf = 6.75 % and Market Risk Premium = MRP = 5 %

Let the portoflio beta be Bp

Therefore, 18 = 6.75 + Bp x 5

11.25 = Bp x 5

Bp = 11.25 / 5 = 2.25

Initial Portfolio Value = $ 20 million , Additional Fund = $ 5 million, and let the average beta of the additional funds be Ba

Initial Portfolio Beta = 2

Therefore, 2.25 = 2 x (20 / 25) + Ba x (5 / 25)

2.25 = 1.6 + 0.2Ba

Ba = (2.25 - 1.6) / 0.2 = 3.25